OpenStudy (anonymous):

Will fan and give medal! Using a directrix of y = -2 and a focus of (2, 6), what quadratic function is created?

OpenStudy (anonymous):

@nickmifsud1223 :)

OpenStudy (anonymous):

I'm sorry but is this multiple choice? xD

OpenStudy (anonymous):

yeah it is @nickmifsud1223 the choices are f(x) = -1/8 (x - 2)^2 - 2 f(x) = 1/16 (x - 2)^2 + 2 f(x) = 1/8 (x - 2)^2 - 2 f(x) = -1/16 (x + 2)^2 - 2

OpenStudy (anonymous):

i guessed and got it right, it was f(x) = 1/16 (x - 2)^2 + 2

OpenStudy (blurbendy):

Do you want an explanation, or are you good?